OpenStudy (achu3p):

Write an equation of the line in standard form that passes through (2, -3) and is perpendicular to -3x+y=2.

OpenStudy (anonymous):

equation of line: y=mx+b, m is the slope, b=y-int since it it is perpendicular to -3x+y=2, it means that the slope of the line is the negative reciprocal of the slope of -3x+y=2.. after you find the slope, find the y-int using the point (2, -3) and y=mx+b equation and plug the slope and the y-int in the equation y=mx+b

OpenStudy (achu3p):

Thank you!